• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 11:16:00 on November16,2009

    Entry number 300322

    keyword=Vacuum is ok. EES temperary swapped a vacuum gauge

    I went to the hall to check the scattering chamber vacuum. The reading was 1.5x10-6. BUt the GUI readback was 2x10-4. Talking to Rick Gonzales from EES group who was trouble shooting the Compton vacuum. He told me that they swapped a vacuum gauge from our target chamber (which caused the analog reading be wrong) to test things out. Later (a few minutes ago) he swapped the gauge back. We are back to normal.